# Synchronization of lasers with complex spatio-temporal dynamics

### v1

@@ II Results
Figure 3(c) shows several regions of positive correlation. The region around $\Delta\nu\approx 0$ GHz stems from synchronization of the $M_{u}(6,1,\pm)$ modes with the $S_{u}(2,1,\pm)$ modes, which are very weak in free-running operation but can be strongly excited by the injection. A closer look shows that there are actually three successive correlation peaks: $S_{u}(2,1,-)$ is excited at $\Delta\nu=-3.5$ GHz, both $S_{u}(2,1,+)$ and $S_{u}(2,1,-)$ are excited at $\Delta\nu=0$ GHz, and $S_{u}(2,1,+)$ is excited at $\Delta\nu=4.14$ GHz as shown in Fig. 3(a). We attribute this behavior to the splitting between the $S_{u}(2,1,-)$ and $S_{u}(2,1,+)$ modes, which seems to be about $\Delta\nu_{O}\approx 7.5$ GHz (see also Appendix A.4): these two modes are excited alone when the $M_{u}(6,1,\pm)$ modes aligns spectrally with them, and their superposition is created when $M_{u}(6,1,\pm)$ is in between the two orientations of $S_{u}(2,1)$. The same behavior is found when the $M_{u}(6,1,\pm)$ modes come close to the $S_{u}(3,1,\pm)$ modes [see Fig. 3(b) and the magnification in Fig. 3(c)]. We associate this scenario to the $M_{u}(6,1,\pm)$ modes lasing simultaneously, which creates an azimuthally uniform intensity distribution that is able to excite both orientations of the slave laser modes equally well.
@@
Applying a low-pass filter improves the correlations to some extent, typically reaching peak values between 20% and 30% with cutoff frequencies around $0.5$ to $1$ GHz [see Fig. 3(c)], but these values are below those found for the first case [see Fig. 2]. We believe this is related to the type the master laser dynamics, which is chaotic with high frequency components up to $20$ GHz [see Fig. 1(c)], in contrast to the slower polarization-hopping dynamics in the first case. Whereas the slow polarization-hopping dynamics creates strong synchronization which is revealed by low-pass filtering, the chaotic dynamics in the second case relies on high-frequency components, so filtering does not help much. It seems that fast chaotic dynamics is harder to synchronize. The absence of inverse synchronization in the second case is explained by the absence of polarization-hopping dynamics to which we attribute the inverse synchronization in the first case. In summary, the second injection experiment confirms the importance of spectral alignment for synchronization [see also Appendix Fig. A12]. However, we also find significant differences between the two cases, demonstrating the diversity of synchronization scenarios and their dependency on the dynamical and spatial properties of the master laser.
III Discussion
Our experiment constitutes the first demonstration of chaos synchronization between two diode lasers which are intrinsically chaotic without optical feedback or external perturbations. Furthermore, it also elucidates the question of synchronization between coupled lasers with complex spatio-temporal dynamics.
@@ III Discussion
Owing to the multimode nature and intrinsic complexity of the VCSELs, the dynamics of the coupled system is very rich, and we observe different scenarios including synchronization of ultrafast chaos, of polarization-hopping dynamics, and inverse synchronization. The measured correlations are generally low of the order of $20\%$, which could be due to the very low transmission through the top mirror of the VCSELs, mismatch in the device structures and the natural strong mode competition. However, correlations up to $90\%$ are observed for low-pass filtered polarization-hopping dynamics. Synchronization is typically observed when one of the dominant modes of the master laser is spectrally aligned with a mode of the slave laser, but does not require a careful matching of their spatial profiles. This flexibility and the multitude of possible synchronization scenarios gives hope that further improvements of the synchronization quality are possible.
@@
Our work opens new perspectives concerning the synchronization of complex lasers. We show that synchronization of the temporal dynamics does not require or imply that other system properties like the spectrum and the spatial profiles become identical as well. Hence a more precise exploration and classification of synchronization scenarios for coupled spatio-temporal systems will be needed. And while we concentrate on the temporal dynamics, further investigations of the spatial dimension appear promising, for example by using different cavity shapes [51, 52, 53] or shaping the injected beam. Tuning the birefringence or introducing controlled anisotropies could help to isolate polarization states and their associated spatial modes. By systematically varying the number and nature of the excited modes, one could assess how the modal complexity influences the robustness, speed, or nature of synchronization. Such control would allow for a more precise exploration of the transition from low-dimensional to high-dimensional chaotic synchronization in spatially extended laser systems.
Finally, there are also perspectives for future applications. The high correlation of the low-pass filtered polarization-hopping dynamics is promising for applications such as private key sharing [54, 55] that typically require both high-dimensional chaos and good synchronization quality, but not very high speeds. Furthermore, broad-area lasers offer the possibility of spatial multiplexing to further increase the processing speed.
@@ Acknowledgements
Competing interests
@@
Figure A4: Experimental setup. Obj1: 40x objective (NA = 0.6); Obj2: 20x objective (NA = 0.5); OI: optical isolator; NDF: neutral density filter; BS: beam splitter; HWP: half-wave plate; FC: fiber coupler; PD: Photodetector; Amp: RF-Amplifier; OSC: oscilloscope; L1 (L2): lens with 200 mm (100 mm) focal length; BD: beam displacer.
Appendix 1 Experimental setup
The experimental setup is presented in Fig. A4 in more detail. The lasers under study are broad-area VCSELs (Frankfurt Laser Company FL85-F1-P1N-AC) with a circular output aperture of $15~{}\mu$m diameter, electrically pumped through a ring contact. The output beam of the master laser is collimated by a 40× microscope objective (Obj1, NA = 0.6). An optical isolator (OI) is placed immediately after the master laser to prevent any undesirable optical feedback that could influence its intrinsic dynamics. A beam splitter (reflexion: 70 / transmission: 30) directs part of the master laser signal to the detection arm (30%), and part of the signal (70%) is injected into the slave laser. In the same way, part of the slave laser signal is collimated and sent to the detection arm.
The detection arm allows to measure the temporal dynamics, the optical spectra, and spatio-spectral images of the two lasers. Two fiber couplers with 20× objectives (Obj2, NA = 0.5) are employed to couple the beams into multimode fibers (Thorlabs M116L02, 50 $\mu m$ core diameter) connected to high-speed photodetectors (Newport 1484-A-50, 22 GHz bandwidth). The RF-signals from the photodetectors are amplified by RF-amplifiers (SHF S126A) yielding a gain of 29 dB over a frequency range spanning from 80 kHz to 25 GHz, and measured by an oscilloscope (Tektronix DPO72340SX, 23 GHz bandwidth). Alternatively, the multimode fibers can be connected to an optical spectrum analyzer (OSA, Anritsu MS9740A, $30$ pm resolution) to examine the spectral composition of the laser emission. Optical isolators are used to prevent reflections from the fiber facets, and the entrance polarizers of the isolators are aligned to transmit only the $u$-polarization of the lasers (unless otherwise noted).
@@ Appendix 1 Experimental setup
The signal of the master laser passes through the collimation objective (Obj1, NA = 0.6), a lens (L1, 200 mm focal length), another lens of the same type and finally the objective (Obj1) in front of the slaver laser. The lenses are aligned such that the output facet of the master laser is imaged onto the output facet of the slave laser with 1x magnification. Furthermore, additional cameras in the far-field planes of the lasers (not shown in Fig. A4) are used to ensure that the injected beam is parallel to the output beam of the slave laser. The optical isolator after the master laser is used to select only its $u$-polarization for transmission to the slave laser, and a half-wave plate is used to align the polarization of the master signal with the $u$-polarization axis of the slave laser for parallel optical injection. A variable neutral density filter allows to change the power of the injected signal. The results presented in this article were obtained for minimal attenuation (maximal injection power).
@@
Appendix 2 Analysis of times series
A Correlation measurement
$C(\Delta t)=\frac{\langle[P_{M}(t)-\langle P_{M}\rangle][P_{S}(t+\Delta t)-\langle P_{S}\rangle]\rangle}{\sqrt{\langle[P_{M}(t)-\langle P_{M}\rangle]^{2}\rangle\langle[P_{S}(t)-\langle P_{S}\rangle]^{2}\rangle}}$ (A1)
[…]
B Filtering of time traces
$H(f)=\frac{1}{\sqrt{1+(f/f_{c})^{2n}}},$ (A2)
where $f_{c}$ is the cutoff frequency. It corresponds to the point where the transfer function drops to -3 dB, marking the boundary between the passband and the stopband. In our case, a filter of order $n=2$ was used, providing a suitable trade-off between high-frequency attenuation and filter stability.
Appendix 3 Properties of free-running lasers
@@ A.1 LI-curves and polarization-switching points
Figure A5: LI-curves of master (a) and slave laser (b) for 19 polarizer angles ranging from 0 to 180. The color corresponds to the polarizer angle, where $u$-polarization (at 90°) is shown in red and $v$-polarization (at 0 and 180) in blue. The dashed green lines indicate the polarization switching points (PSPs).
To identify the principal polarization axes of the BA-VCSELs used in this work, a polarizer in a motorized rotation stage was employed to measure the optical power transmitted through the polarizer as function of its orientation for both lasers. Using a power meter, measurements were taken every 10 over the full current range of 2 to 10 mA as shown in Fig. A5. The resulting profiles exhibit two distinct extrema in emitted power at 0 (or equivalently 180) and 90, revealing the two orthogonal polarization axes. Depending on the pump current and the temperature of the laser, the dominant polarization switches between these two orientations at specific polarization switching points (PSPs), while the $u$- and $v$-polarization axes themselves remain fixed. In most cases, the 90 axis (referred to as $u$-polarization) dominates, while the 0 / 180 axis is referred to as $v$-polarization. See Ref. [38] for more details. It should be noted that we define the polarizer angle with respect to the polarization axes of the VCSELs as a convenient reference frame to facilitate visualization and comparison. These axes generally do not coincide with the horizontal and vertical axes of the laboratory frame.
@@ A.2 Optical spectrum and birefringence
Figure A6: Optical spectra of the master laser at a pump current of $I_{M}=6.5$ mA and a temperature of $T_{M}=20^{\circ}$, measured along the $u$- (red) and $v$-polarization (blue) axes.
Measurements of the optical spectra provide important information on the properties of the lasers under study. First, the wavelengths of the individual modes are clearly identified, with the fundamental mode appearing at the longest wavelength (right side of the spectrum), and higher-order transverse modes at shorter wavelengths (as discussed in more detail in Ref. [38]). Figure A6 shows the optical spectra of the master laser measured along the two dominant polarization axes, $u$ and $v$, for a fixed pump current of $I_{M}=6.5$ mA. This measurement reveals that the optical power is distributed differently among the transverse modes depending on the polarization, and also highlights a spectral shift between the two polarization components. We assign the label ”red” (”blue”) to the $u$- ($v$-) polarization whose spectrum is at longer (shorter) wavelengths.
@@ A.3 Polarization-hopping dynamics
Figure A7: (a) Time traces of the master laser at $I_{M}$= 4.95 mA, low-pass filtered at a cutoff of 0.1 GHz, for the two orthogonal polarization axes $u$ (red) and $v$ (blue). (b) Mean dwell time of the time traces in one polarization state as function of the pump current.
The VCSELs under study exhibit several polarization switching points (PSPs) at which the dominant polarization changes (see Fig. A5), such as in the case of the master laser at 3.6 mA, as well as points at which the power in the two polarization states becomes almost equal without switching, such as the master laser at 6.2 mA. These points are sometimes accompanied by a dynamical regime known as polarization hopping [47]. As illustrated in Fig. A7(a), this regime is characterized by intensity fluctuations in one polarization ($u$), with an anti-correlated response in the orthogonal polarization ($v$). Due to the relatively long timescales of this dynamics, the associated intensity fluctuations contribute significantly to the low-frequency components of the RF spectrum, predominantly below $500$ MHz.
This behavior has been reported in various VCSEL systems [44, 45, 46, 47], and the spontaneous polarization switching can be driven by noise or by deterministic nonlinear dynamics. The temporal behavior within this regime is characterized by the dwell time, defined as the average duration the system remains in one polarization state before switching. Figure A7(b) presents the measured dwell time as a function of pump current, showing an exponential increase with increasing current. This trend is characteristic of deterministic polarization chaos [46].
@@ A.4 Orientation of transverse modes
Figure A8: Spatio-spectral image of the free-running master laser in $u-$polarization for $I_{M}$= 9.8 mA with the modes $M_{u}(7,1,+)$ and $M_{u}(7,1,-)$ highlighted in the inset in blue. They are offset laterally due othe the non-vanishing orientation splitting $\Delta\nu_{O}\approx 11$ GHz.
Due to the circular geometry of the VCSEL cavity, all transverse modes $(m,n)$ with $m>0$ exist in two different orientations, designated by $+$ (--) if they are (anti-) symmetric with respect to a chosen axis. In an ideally symmetric cavity, the modes with $+$ and -- orientations are degenerate.
Figure 1(b2) presents polarization-resolved spatio-spectral measurements of the master laser under free-running conditions for $I_{M}$= 6.5 mA. Several key observation can be made: first, we typically observe modes with the same transverse pattern in both polarization states, however, there is a clear spectral shift between them due to the birefringence as discussed above. Second, the intensities in the two polarizations are generally not the same (see Fig. A6). Third, for a given $(m,n)$, the $u$- and $v$-polarizations usually exhibit different transverse orientations, with a -- symmetry in $u$-polarization corresponding to a $+$ symmetry in $v$-polarization, which is attributed to gain competition [41, 42, 43]. In short, usually each transverse mode lases in both polarizations but with different spatial orientations and different intensities. However, in some cases, such as mode $M_{u}(6,1)$ in Fig. 3, the measured spatial pattern shows that both orientations lase in the same polarization state (under free-running conditions).
Figure A8 shows the spatio-spectral image of the master laser at 9.8 mA. In this example, the mode $(7,1)$ lases in both orientations in $u$-polarization, however, we see that the $M_{u}(7,1,+)$ and the $M_{u}(7,1,-)$ mode are not degenerate as expected for a perfectly symmetric system, but spectrally shifted. We designate this frequency separation between the $+$ and -- orientations of a given transverse mode within the same polarization as $\Delta\nu_{O}$. This splitting is usually below the resolution of standard optical spectrum analyzers or imaging spectrometers, making it difficult to measure directly. In the case of Fig. A8, it is quite large with $\Delta\nu_{O}\approx 11$ GHz. In other instances, like the mode $M_{u}(6,1)$ in Fig. 3, it is close to 00.

### v2

@@ Results
Figure 3(c) shows several regions of positive correlation. The region around $\Delta\nu\approx 0$ GHz stems from synchronization of the $M_{u}(6,1,\pm)$ modes with the $S_{u}(2,1,\pm)$ modes, which are very weak in free-running operation but can be strongly excited by the injection. A closer look shows that there are actually three successive correlation peaks: $S_{u}(2,1,-)$ is excited at $\Delta\nu=-3.5$ GHz, both $S_{u}(2,1,+)$ and $S_{u}(2,1,-)$ are excited at $\Delta\nu=0$ GHz, and $S_{u}(2,1,+)$ is excited at $\Delta\nu=4.14$ GHz as shown in Fig. 3(a). We attribute this behavior to the splitting between the $S_{u}(2,1,-)$ and $S_{u}(2,1,+)$ modes, which seems to be about $\Delta\nu_{O}\approx 7.5$ GHz (see also Supplementary Section A.4): these two modes are excited alone when the $M_{u}(6,1,\pm)$ modes aligns spectrally with them, and their superposition is created when $M_{u}(6,1,\pm)$ is in between the two orientations of $S_{u}(2,1)$. The same behavior is found when the $M_{u}(6,1,\pm)$ modes come close to the $S_{u}(3,1,\pm)$ modes [see Fig. 3(b) and the magnification in Fig. 3(c)]. We associate this scenario to the $M_{u}(6,1,\pm)$ modes lasing simultaneously, which creates an azimuthally uniform intensity distribution that is able to excite both orientations of the slave laser modes equally well.
@@
Applying a low-pass filter improves the correlations to some extent, typically reaching peak values between 20% and 30% with cutoff frequencies around $0.5$ to $1$ GHz [see Fig. 3(c)], but these values are below those found for the first case [see Fig. 2]. We believe this is related to the type the master laser dynamics, which is chaotic with high frequency components up to $20$ GHz [see Fig. 1(c)], in contrast to the slower polarization-hopping dynamics in the first case. Whereas the slow polarization-hopping dynamics creates strong synchronization which is revealed by low-pass filtering, the chaotic dynamics in the second case relies on high-frequency components, so filtering does not help much. It seems that fast chaotic dynamics is harder to synchronize. The absence of inverse synchronization in the second case is explained by the absence of polarization-hopping dynamics to which we attribute the inverse synchronization in the first case. In summary, the second injection experiment confirms the importance of spectral alignment for synchronization [see also Supplementary Fig. S9]. However, we also find significant differences between the two cases, demonstrating the diversity of synchronization scenarios and their dependency on the dynamical and spatial properties of the master laser.
Discussion
Our experiment constitutes a lab-scale demonstration of synchronization between systems with complex spatio-temporal dynamics. By exploiting the intrinsic and controllable chaotic dynamics of broad-area VCSELs, it also constitutes the simplest and most practical architecture so far for multiple-user physical-layer secure communication schemes at high bit rates. Such systems would benefit from the highly complex and fast fluctuations of the laser dynamics and the spatial aspect of the dynamics involving a large number of laser modes. Moreover, the high-dimensional spatio-temporal dynamics of broad-area VCSELs could also be exploited in future photonic reservoir-computing architectures [53].
@@ Discussion
Owing to the multimode nature and intrinsic complexity of the VCSELs, the dynamics of the coupled system is very rich, and we observe different scenarios including synchronization of ultrafast chaos, polarization-hopping dynamics, and inverse synchronization. The measured correlations are generally low of the order of $20\%$, which could be due to the very low transmission through the top mirror of the VCSELs, mismatch in the device structures and the intrinsic mode competition. However, correlations up to $90\%$ are observed for low-pass filtered polarization-hopping dynamics. Synchronization is typically observed when one of the dominant modes of the master laser is spectrally aligned with a mode of the slave laser, but does not require a careful matching of their spatial profiles. This flexibility and the multitude of possible synchronization scenarios gives hope that further improvements of the synchronization quality are possible. Still, the high correlation of the low-pass filtered polarization-hopping dynamics is promising for applications such as private key sharing [4, 27] that typically require both high-dimensional chaos and good synchronization quality, but not very high speeds.
@@
Our work opens new perspectives concerning the synchronization of complex lasers and beyond. We show that synchronization of the temporal dynamics does not require or imply that other system properties like the spectrum and the spatial profiles become identical as well. Hence a more precise exploration and classification of synchronization scenarios for coupled spatio-temporal systems will be needed, with implications in other scientific fields. Beyond the temporal dynamics, further investigations of the spatial dimension appear promising, for example by using different cavity shapes [14, 28, 3] or shaping the injected beam. Tuning the birefringence or introducing controlled anisotropies could help to isolate polarization states and their associated spatial modes. By systematically varying the number and nature of the excited modes, one could assess how the modal complexity influences the robustness, speed, or nature of synchronization. Such control would allow for a more precise exploration of the transition from low-dimensional to high-dimensional chaotic synchronization in spatially extended laser systems.
Materials and Methods
Correlation measurement
$C(\Delta t)=\frac{\langle[P_{M}(t)-\langle P_{M}\rangle][P_{S}(t+\Delta t)-\langle P_{S}\rangle]\rangle}{\sqrt{\langle[P_{M}(t)-\langle P_{M}\rangle]^{2}\rangle\langle[P_{S}(t)-\langle P_{S}\rangle]^{2}\rangle}}$ (1)
where $P_{M}(t)$ and $P_{S}(t)$ are the output powers (i.e., time traces) of the master and slave lasers, respectively, that are measured with the high-speed photodetectors. The time average is denoted by $\langle\rangle$, and $\Delta t$ is a time shift. The correlation is computed as function of $\Delta t$ for each detuning $\Delta\nu$. Figs. 2(b) and 3(c), however, only show the correlation at the $\Delta t$ for which it is maximal. This time shift of $\Delta t\approx 3.78$ ns corresponds to the path-length difference between the master and slave signals, and the maximum of the correlation is indeed obtained at this delay, confirming that zero-lag synchronization is computed and observed. For detunings where the master and slave lasers are not synchronized, the correlation is theoretically expected to vanish. However, a persistent residual correlation remains observable in our system, fluctuating around $C=3\%$ for the unfiltered data. This non-vanishing baseline arises from the reflection of the master laser from the top mirror of the slave laser. We verified that even with the slave laser switched off, a small fraction of the master signal is reflected back, producing a weak but measurable correlation. This weak correlation also appears at a time lag $\Delta t$ corresponding to the propagation delay along the signal pathways.
Filtering of time traces
$H(f)=\frac{1}{\sqrt{1+(f/f_{c})^{2n}}}$ (2)
where $f_{c}$ is the cutoff frequency. It corresponds to the point where the transfer function drops to -3 dB, marking the boundary between the passband and the stopband. In our case, a filter of order $n=2$ was used, providing a suitable trade-off between high-frequency attenuation and filter stability. We verified that the type and order of the low-pass filter (Butterworth, zero-phase, or linear-phase FIR) do not affect the synchronization coefficient, indicating that filtering does not distort the phase or delay in a way that impacts our results.
@@ Acknowledgements
Data availability
Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
Conflict of interest
@@
Supplementary Information for ”Synchronization of complex spatio-temporal dynamics with lasers”
adf
Jules Mercadier∗,1,2, Stefan Bittner1,2 and Marc Sciamanna∗,1,2
1Université de Lorraine, CentraleSupélec, LMOPS F-57000 Metz, France
2Chaire Photonique, LMOPS, CentraleSupélec, 57070 Metz, France
Corresponding authors : jules.mercadier@centralesupelec.fr; marc.sciamanna@centralesupelec.fr
Figure S1: Experimental setup. Obj1: 40x objective (NA = 0.6); Obj2: 20x objective (NA = 0.5); OI: optical isolator; NDF: neutral density filter; BS: beam splitter; HWP: half-wave plate; FC: fiber coupler; PD: Photodetector; Amp: RF-Amplifier; OSC: oscilloscope; L1 (L2): lens with 200 mm (100 mm) focal length; BD: beam displacer.
Supplementary Section 1 Experimental setup
[…]
@@
The signal of the master laser passes through the collimation objective (Obj1, NA = 0.6), a lens (L1, 200 mm focal length), another lens of the same type and finally the objective (Obj1) in front of the slaver laser. The lenses are aligned such that the output facet of the master laser is imaged onto the output facet of the slave laser with 1x magnification. Furthermore, additional cameras in the far-field planes of the lasers (not shown in Fig. S1) are used to ensure that the injected beam is parallel to the output beam of the slave laser. The optical isolator after the master laser is used to select only its $u$-polarization for transmission to the slave laser, and a half-wave plate is used to align the polarization of the master signal with the $u$-polarization axis of the slave laser for parallel optical injection. A variable neutral density filter allows to change the power of the injected signal. The results presented in this article were obtained for minimal attenuation (maximal injection power).
@@
Supplementary Section 2 Properties of free-running lasers
@@ A.1 LI-curves and polarization-switching points
Figure S2: LI-curves of master (a) and slave laser (b) for 19 polarizer angles ranging from 0 to 180. The color corresponds to the polarizer angle, where $u$-polarization (at 90°) is shown in red and $v$-polarization (at 0 and 180) in blue. The dashed green lines indicate the polarization switching points (PSPs).
To identify the principal polarization axes of the BA-VCSELs used in this work, a polarizer in a motorized rotation stage was employed to measure the optical power transmitted through the polarizer as function of its orientation for both lasers. Using a power meter, measurements were taken every 10 over the full current range of 2 to 10 mA as shown in Fig. S2. The resulting profiles exhibit two distinct extrema in emitted power at 0 (or equivalently 180) and 90, revealing the two orthogonal polarization axes. Depending on the pump current and the temperature of the laser, the dominant polarization switches between these two orientations at specific PSPs, while the $u$- and $v$-polarization axes themselves remain fixed. In most cases, the 90 axis (referred to as $u$-polarization) dominates, while the 0 / 180 axis is referred to as $v$-polarization. See Ref. [S1] for more details. It should be noted that we define the polarizer angle with respect to the polarization axes of the VCSELs as a convenient reference frame to facilitate visualization and comparison. These axes generally do not coincide with the horizontal and vertical axes of the laboratory frame.
@@ A.2 Optical spectrum and birefringence
Figure S3: Optical spectra of the master laser at a pump current of $I_{M}=6.5$ mA and a temperature of $T_{M}=20^{\circ}$C, measured along the $u$- (red) and $v$-polarization (blue) axes.
Measurements of the optical spectra provide important information on the properties of the lasers under study. First, the wavelengths of the individual modes are clearly identified, with the fundamental mode appearing at the longest wavelength (right side of the spectrum), and higher-order transverse modes at shorter wavelengths (as discussed in more detail in Ref. [S1]). Figure S3 shows the optical spectra of the master laser measured along the two dominant polarization axes, $u$ and $v$, for a fixed pump current of $I_{M}=6.5$ mA. This measurement reveals that the optical power is distributed differently among the transverse modes depending on the polarization, and also highlights a spectral shift between the two polarization components. We assign the label ”red” (”blue”) to the $u$- ($v$-) polarization whose spectrum is at longer (shorter) wavelengths.
@@ A.3 Polarization-hopping dynamics
Figure S4: (a) Time traces of the master laser at $I_{M}$= 4.95 mA, low-pass filtered at a cutoff of 0.1 GHz, for the two orthogonal polarization axes $u$ (red) and $v$ (blue). (b) Mean dwell time of the time traces in one polarization state as function of the pump current.
The VCSELs under study exhibit several polarization switching points (PSPs) at which the dominant polarization changes (see Fig. S2), such as in the case of the master laser at 3.6 mA, as well as points at which the power in the two polarization states becomes almost equal without switching, such as the master laser at 6.2 mA. These points are sometimes accompanied by a dynamical regime known as polarization hopping. As illustrated in Fig. S4(a), this regime is characterized by intensity fluctuations in one polarization ($u$), with an anti-correlated response in the orthogonal polarization ($v$). Due to the relatively long timescales of this dynamics, the associated intensity fluctuations contribute significantly to the low-frequency components of the RF spectrum, predominantly below $500$ MHz.
This behavior has been reported in various VCSEL systems [S2-S4], and the spontaneous polarization switching can be driven by noise or by deterministic nonlinear dynamics. The temporal behavior within this regime is characterized by the dwell time, defined as the average duration the system remains in one polarization state before switching. Figure S4(b) presents the measured dwell time as a function of pump current, showing an exponential increase with increasing current. This trend is characteristic of deterministic polarization chaos [S4].
@@ A.4 Orientation of transverse modes
Figure S5: Spatio-spectral image of the free-running master laser in $u$-polarization for $I_{M}$= 9.8 mA with the modes $M_{u}(7,1,+)$ and $M_{u}(7,1,-)$ highlighted in the inset in blue. They are offset laterally due othe the non-vanishing orientation splitting $\Delta\nu_{O}\approx 11$ GHz.
Due to the circular geometry of the VCSEL cavity, all transverse modes $(m,n)$ with $m>0$ exist in two different orientations, designated by ++ (-) if they are (anti-) symmetric with respect to a chosen axis. In an ideally symmetric cavity, the modes with ++ and - orientations are degenerate.
Figure 1(b2) presents polarization-resolved spatio-spectral measurements of the master laser under free-running conditions for $I_{M}$= 6.5 mA. Several key observation can be made: first, we typically observe modes with the same transverse pattern in both polarization states, however, there is a clear spectral shift between them due to the birefringence as discussed above. Second, the intensities in the two polarizations are generally not the same (see Fig. S3). Third, for a given $(m,n)$, the $u$- and $v$-polarizations usually exhibit different transverse orientations, with a - symmetry in $u$-polarization corresponding to a ++ symmetry in $v$-polarization, which is attributed to gain competition [S5,S6]. In short, usually each transverse mode lases in both polarizations but with different spatial orientations and different intensities. However, in some cases, such as mode $M_{u}(6,1)$ in Fig. 3, the measured spatial pattern shows that both orientations lase in the same polarization state (under free-running conditions).
Figure S5 shows the spatio-spectral image of the master laser at 9.8 mA. In this example, the mode $(7,1)$ lases in both orientations in $u$-polarization, however, we see that the $M_{u}(7,1,+)$ and the $M_{u}(7,1,-)$ mode are not degenerate as expected for a perfectly symmetric system, but spectrally shifted. We designate this frequency separation between the ++ and - orientations of a given transverse mode within the same polarization as $\Delta\nu_{O}$. This splitting is usually below the resolution of standard optical spectrum analyzers or imaging spectrometers, making it difficult to measure directly. In the case of Fig. S5, it is quite large with $\Delta\nu_{O}\approx 11$ GHz. In other instances, like the mode $M_{u}(6,1)$ in Fig. 3, it is close to 0.